OpenStudy (anonymous):

What is the correct solution to the equation? 5(x-5) = 15 a: -50 b: -2 c: 8 d: 80

OpenStudy (ageta):

Let's solve your equation step-by-step. 5(x−5)=15 Step 1: Simplify both sides of the equation. 5x−25=15 Step 2: Add 25 to both sides. 5x−25+25=15+25 5x=40 Step 3: Divide both sides by 5.

OpenStudy (ageta):

Answer: x=8

OpenStudy (ageta):

does that help?

OpenStudy (anonymous):

its better to draw it out

OpenStudy (anonymous):